40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Nov 12 08:21:22 2021, Aidan, Summary, 2um Photodiodes, Results from JPL PD: A1-Test3 A1_test3_nominal_QE.pdfA1_test3_darkcurrent.pdfA1_test3_noise_spectra_rev.pdf
    Reply  Fri Nov 12 10:54:55 2021, rana, Summary, 2um Photodiodes, Results from JPL PD: A1-Test3 
Message ID: 2692     Entry time: Fri Nov 12 08:21:22 2021     Reply to this: 2693
Author: Aidan 
Type: Summary 
Category: 2um Photodiodes 
Subject: Results from JPL PD: A1-Test3 

[Aidan]

Here is the analyzed data from Test 3 of the A1 JPL PD.

Note about the QE measurement: 

  • HOM beam QE was performed during the warm up phase. Collimating lens was fixed and beam pointing was optimized at 100mA before each measurement. Likely that not all power was on PD but that distribution was constant throughout measurement. Therefore, good proxy for shape of QE response.
  • Manual QE was performed with 25mA current, optimized collimating lens position (and thus the beam size on the PD). The data corresponds to 8.0mm between the lens mount and fiber mount. The beam pointing was optimized before each measurement at 25mA.
  • QE projection scales the "HOM beam QE" result to the manual QE measurements to project out expected QE performance vs temperature

Dark current is the output from the Keithley scan - the vertical scale is correct in Amps [ignore question mark]

Dark noise spectra are included for different bias levels and at different temperatures. Stll need to add ADC noise floor for these plots.

Notes from Test 3

~/JPL_PD/data/A1_test3/README

6-Oct-2021: done with cryocooler off and temperature increasing

PREAMP GAIN = 1E3

SR560 gain = 500

LD temp set point = 20.2kOhm

 

 

 

Attachment 1: A1_test3_nominal_QE.pdf  22 kB  | Hide | Hide all
A1_test3_nominal_QE.pdf
Attachment 2: A1_test3_darkcurrent.pdf  18 kB  | Hide | Hide all
A1_test3_darkcurrent.pdf
Attachment 3: A1_test3_noise_spectra_rev.pdf  1.372 MB  | Hide | Hide all
A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf
ELOG V3.1.3-